Find the first five terms of each geometric sequence described.
1. a sub1=-1, r=-3
2. a sub1=-1/3, r=2

Answer by edjones(8007)

1)
a[1]=-1, r=-3
a[2]=-1*-3=3
a[3]=-1(-3)^2=-9
a[4]=-1(-3)^3=27
a[4]=-1(-3)^4=-81
.
Ed